ten years ago a man was as old as his wife is now. The man is now twice as old as his wife was then. How old are they both now?
Answer by mananth(16949)

ten years ago a man was as old as his wife is now. The man is now twice as old as his wife was then. How old are they both now?
let the wife's age be x now
Let the husband's age be y now
y-10 =x
y=2(x-10)
y-x=10
y-2x=-20
y-x -y+2x=-20-10
-x=-30
x= 30 wife's age now
y-10=x
y-10=30
y=40 the husband's age now